Using Mathematica:(2x – 3y = 4 4. Consider the system of equations: 1-2 +1.5y = 3 (a) Graph the two lines corresponding to this system, and use
Answer #1

4. а) The system of equation 2x – 3 -4 1, -x+15 °3 ) іх 2х-3) =41 -x+2ə - з 4 25 – 3 =4 7. – 27 + 3 = 6 ч 2x – 3 = 4 + 0 2 -Here the two lines are parallel. so there are no solution . Using the graph we decid the system . has no solution. CE .Read Carefully.